OpenStudy (countrygirl1431):

Please Help!!!!!!! If 6:2=8:x, then 2x = 14 2x = 48 6x = 10 6x = 16

OpenStudy (texaschic101):

ratios can be written as fractions 6:2 can be written as 6/2 and 8:x can be written as 8/x. So we have : 6/2 = 8/x -- this is a proportion so we cross multiply (6)(x) = (8)(2) 6x = 16 any questions ?

OpenStudy (countrygirl1431):

how did u get 16?

OpenStudy (anonymous):



OpenStudy (texaschic101):

8 x 2 = 16

OpenStudy (countrygirl1431):

oh ok i got it now thanks

OpenStudy (texaschic101):

oh...I crossed multiplied because of the proportion

OpenStudy (countrygirl1431):

i got it thanks

OpenStudy (texaschic101):

YW :)